OpenStudy (anonymous):

2x^4+3=12 but x must be a decimal

OpenStudy (tkhunny):

Do you have a calculator with a square root button?

OpenStudy (anonymous):

ya

OpenStudy (tkhunny):

That's about all you need. 2x^4 + 3 = 12 Subtract 3 2x^4 = 9 Divide by 2 x^4 = 9/2 \(x^{2} = \pm 3/\sqrt{2}\) Now what?

OpenStudy (anonymous):

I sqrt \[\pm3\sqrt{2}\]

OpenStudy (anonymous):

Wait no u cant have a sqrt as a denominator so i have to mutiply both the numerator and denominator by sqrt2

OpenStudy (tkhunny):

I don't normally care about that, but we're not done! That's only x^2. You need 'x'. I'm assuming we're staing in the Real Numbers. If we have x^2 that CANNOT be negative, so we're down to \(x^{2} = \dfrac{3}{\sqrt{2}}\). Can you find x?

OpenStudy (anonymous):

is it \[\pm1.46\] because i have to round to two decimal places

OpenStudy (tkhunny):

You keep making large jumps. Try to figure out how to go one step at a time. We have \(x^{2} = \dfrac{3}{\sqrt{2}}\). This gives \(x = \pm\dfrac{\sqrt{3}}{\sqrt[4]{2}}\).

OpenStudy (tkhunny):

This may just be me, but I wuold suggest ALWAYS giving an exact answer and supply the less exact answer if required.

OpenStudy (anonymous):

okay